# F-95000 Cergy-Pontoise, France

### v1

@@ 4.2 Repairs of Star-Tables
Let $T$ be a star-table over universe $U$. For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{k}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$. Then, for every $R$ in ${\sf Rep}(T)$, there exists $i_{0}$ in $\{1,\ldots,k\}$ such that $xa_{i_{0}}$ is in ${\sf True}(\mathcal{R})$.
Proof. It is easy to show that the proposition holds for $k=1$, since in this case $xa_{1}$ is in ${\sf Cons}(\mathcal{T})$, and thus in ${\sf True}(\mathcal{R})$ for every $R$ in ${\sf Rep}(T)$. We now assume that $k>1$, which implies that for every $\sigma$ in $m\_Chase(T)$ such that $XA\subseteq sch(\sigma)$ and $x$ is in ${\sf tuples}(\sigma(X))$, ${\sf tuples}(\sigma(XA))=(xa_{1},\ldots,xa_{k})$ with $k>1$. Moreover, given the form of functional dependencies in $FD$, either $(a)$ $X=\mathbf{K}$ and $A$ is in $\mathbf{M}$, or $(b)$ there exists $i_{0}$ in $\{1,\ldots,n\}$ such that $X=K_{i_{0}}$ and $A$ is in $sch^{*}(D_{i_{0}})$.
Let $R$ be in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no $xa_{l}$ from $\{xa_{1},\ldots,xa_{k}\}$, and let $l_{0}$ in $\{1,\ldots,k\}$. Denoting by $R^{\prime}$ the table $R\cup\{xa_{l_{0}}\}$ we show that in either case $(a)$ or $(b)$ above, $R^{\prime}\models FD$. To this end, we first notice that, since for every $k$ over $\mathbf{K}$ and for any $k_{i}$ over $K_{i}$, $k$ and $k_{i}$ are in ${\sf Cons}(\mathcal{T})$, the tuple $x$ is in ${\sf True}(\mathcal{R})$. Hence, ${\sf True}(\mathcal{R})$ contains tuples whose $X$-value is $x$ and for every such tuple, the $A$-value is missing. The cases $(a)$ and $(b)$ are then handled as follows when computing $m\_Chase(R^{\prime})$: every $t$ in ${\sf True}(\mathcal{R}^{\prime})$ such that $t.X=x$ is considered and its $A$-value is set to $a_{l_{0}}$. Denoting by $\Sigma$ the resulting m-table and let $X^{\prime}\to A^{\prime}$ be a dependency in $FD$, and let $\sigma_{1}$ and $\sigma_{2}$ be two m-tuples in $\Sigma$ such that $\sigma_{1}(X^{\prime})=\sigma_{2}(X^{\prime})$ where $X^{\prime}$ is either $\mathbf{K}$ or $K_{j}$.
-- If $X^{\prime}=X$ then by definition of $\Sigma$, $\sigma_{1}(A)=\sigma_{2}(A)=(a_{l_{0}})$. Thus $\Sigma$ is not changed.
-- If $X^{\prime}=K_{j}$ and $X=\mathbf{K}$, since as above we have $A^{\prime}\neq A$, $m\_Chase$ does not change $\Sigma$.
-- If $X^{\prime}=\mathbf{K}$ and $X=K_{i}$, then in ${\sf True}(\mathcal{R}^{\prime})$, we have $\sigma_{1}(X)=\sigma_{2}(X)=x$, and $A$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$. By definition of $\Sigma$, $\sigma_{1}(A)$ and $\sigma_{2}(A)$ are set to $(a_{l_{0}})$, and then here again, $m\_Chase$ does not change $\Sigma$.
We thus obtain that $\Sigma=m\_Chase(\mathcal{R}^{\prime})$ and that $R^{\prime}\models FD$, because $R\models FD$ and no conflicts occur in $\Sigma$. Moreover, as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}^{\prime})$. It follows from Definition 2 that $R$ can not be in ${\sf Rep}(T)$ since the inclusion ${\sf True}(\mathcal{R})\subset{\sf True}(\mathcal{R}^{\prime})$ is strict (because $xa_{l_{0}}$ is in ${\sf True}(\mathcal{R}^{\prime})$ but not in ${\sf True}(\mathcal{R})$). As this is a contradiction with our hypothesis that $R$ is in ${\sf Rep}(T)$, the proof is complete. $\Box$
@@
We now show how to characterize all repairs of a star-table $T$, based on $m\_Chase(T)$. To this end, we define the following process (P):
@@
Step P1. For every $i$ in $\{1,\ldots,n\}$ and every $A_{i}^{j}$ in $sch^{*}(D_{i})$, if $k_{i}$ is a $K_{i}$-value occurring in $m\_Chase(T)$, we choose one $A_{i}^{j}$-value in the set $\sigma(A_{i}^{j})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$. Denoting by $\varphi_{i}^{j}(k_{i})$ this value, we notice that $\varphi_{i}^{j}(k_{i})$ is well defined because, thanks to $K_{i}\to A_{i}^{j}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$ have the same value over $A_{i}^{j}$.
@@
Step P2. For every $l$ in $\{1,\ldots,p\}$, if $k$ is a $\mathbf{K}$-value occurring in $m\_Chase(T)$, we choose one $M_{l}$-value in the set $\sigma(M_{l})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$. Denoting by $\varphi_{l}(k)$ this value, we notice that, as above, $\varphi_{l}(k)$ is well defined because, thanks to $\mathbf{K}\to M_{l}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$ have the same value over $M_{l}$.
@@
Step P3. Denoting by $\varphi$ the set of all $\varphi_{i}^{j}(k_{i})$ and all $\varphi_{l}(k)$ defined above, for every $\sigma$ in $m\_Chase(T)$, let $t_{\varphi}(\sigma)$ be the tuple such that $sch(t_{\varphi}(\sigma))=sch(\sigma)$, defined by:
@@
  • If $\mathbf{K}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$, for every $l=1,\ldots,p$ such that $M_{l}\in sch(\sigma)$, let $t_{\varphi}(\sigma).M_{l}=\varphi_{l}(k)$.
@@
In the context of Example 3, $T$ has four repairs as shown in Figure 3. It can be seen from the m-table $m\_Chase(T)$ in Figure 2 that, in each of these four tables:
  • the first two rows are generated using the first m-tuple in $m\_Chase(T)$, namely
@@
  • the next five rows are generated using the second m-tuple in $m\_Chase(T)$, namely
@@
  • the last two rows are generated using respectively the last two m-tuples in $m\_Chase(T)$, namely $(k^{\prime}_{1})(k^{\prime\prime}_{2})(a_{1})(m_{1})$ and $(k^{\prime\prime}_{1})(a^{\prime}_{1})(a^{\prime}_{2})$.
@@
  • Step P2. The only measure attribute is here $M_{1}$ and the $\mathbf{K}$-values to be considered are $k_{1}k_{2}$, $k_{1}k^{\prime}_{2}$ and $k^{\prime}_{1}k^{\prime\prime}_{2}$ which are respectively associated with $M_{1}$-values $m_{1}$, $m^{\prime\prime}_{1}$ and $m_{1}$.
  • Step P3. Considering every m-tuple $\sigma$ of the table $m\_Chase(T)$ shown in Figure 2, we obtain the following four tuples: $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$ and $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
We obtain that $R_{\varphi}$ contains the four tuples above along with the tuples in ${\sf Cons}(\mathcal{T})$ that have been characterized in Example 3. Moreover, inspecting thoroughly the table $R_{1}$ and the set $R_{\varphi}$ shows that the two sets yield the same set of true tuples, corresponding to the same repair of $T$. $\Box$
@@ 5 Consistent Query Answering
    Standard queries or simply queries, to be defined in the next section. These queries are simple projection-selection queries of the form select $X$ from $T$ where $\Gamma$, where $X$ is a list of attributes and $\Gamma$ an optional selection condition.
@@
    Analytic queries with no group-by statement, defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$, where $X$ and $\Gamma$ are as above, and where $aggr$ is an aggregation operator and $M_{i}$ is a measure attribute.
@@
    Analytic queries with a group-by statement, also defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above.
@@
    Analytic queries with a group-by-having statement, dealt with in Section 6.4. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above, and where $\Upsilon$ is a boolean expression involving aggregates.
@@ 5.1 Standard Queries
The set of all attributes occurring in $\Gamma$ is called the schema of $\Gamma$, denoted by $sch(\Gamma)$; and the attribute set $X\cup sch(\Gamma)$ is called the schema of $Q$, denoted by $sch(Q)$.
@@
Given a selection condition $\Gamma$, we denote by $Sat(\Gamma)$ the set of all tuples in $\mathcal{T}(sch(\Gamma))$ satisfying $\Gamma$, as defined below:
@@
Moreover, based on Proposition 1, it is shown in LauS25 that the two bounds ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ can be easily computed by means of one scan of the m-table $m\_Chase(T)$. It turns out that approximating the consistent answer to any query in our approach is polynomial. For example, in the context of Example 3,
@@ 5.2 Consistent Query Answering and Star Schemas
As mentioned in our introductory section, we can compute efficiently exact consistent answers (instead of bounds of exact consistent anwers as in LauS25 ) if the selection condition meets two requirements. The first requirement is that satisfaction of the selection condition can be tested attribute by attribute, as stated in the following definition.
@@
For example, $\Gamma=(A_{1}\leq 0)\vee((A_{1}\geq 4)\wedge((A_{2}=a))$ can be written as $((A_{1}\leq 0)\vee(A_{1}\geq 4))\wedge((A_{1}\leq 0)\vee(A_{2}=a))$. Thus, $\Gamma$ is not independent. On the other hand, $\Gamma^{\prime}=((A_{1}\leq 4)\vee(A_{1}\geq 0))\wedge(A_{2}=a)$ is obviously independent with $\Gamma^{\prime}(A_{1})=(A_{1}\geq 4)\vee(A_{1}\leq 0)$ and $\Gamma^{\prime}(A_{2})=(A_{2}=a)$, and we have $Sat(\Gamma^{\prime})=\{(a_{1},a_{2})\in adom(A_{1})\times adom(A_{2})~{}|~{}(a_{1}\leq 0\vee a_{1}\geq 4)\wedge(a_{2}=a)\}$.
@@
Another restriction that we consider on queries is that key attributes in $\mathbf{K}$ are not allowed in selection conditions. We point out that this syntactical restriction on queries does not rule out queries of practical interest, because in most relevant queries, selections are expressed using meaningful attributes rather than numerical abstract key-values. Notice in this respect that key attributes are mainly used in selection conditions to express join conditions, which are not explicit in our chase-based approach.
  • From now on, it is assumed, even when not explicitly mentioned, that selection conditions are independent and involve no key attributes from $\mathbf{K}$.
Actually, what the above assumption implies is that, in a star-table, given a query $Q:{\tt select}$ $X$ from $T$ where $\Gamma$, for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$ holds, we have $Y\subseteq X\setminus sch(\Gamma)$.
@@
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
@@
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
  3. 3.
    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
Proof. See Appendix B. $\Box$
An important consequence of Proposition 8 is that the data complexity of consistent query answering to projection-selection-join queries on star schemes is polynomial, under the restriction that the selection condition is independent and involves no keys. The following corollary shows that the third item in Proposition 8 can be more simply stated when $\mathbf{K}$ is a subset of $X$.
@@
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. If $\mathbf{K}\subseteq X$, ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
@@
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
@@
Proof. As for every $A$ in $U$, there exists $Y\to A$ in $FD$ with $Y\subseteq\mathbf{K}$, if $\mathbf{K}\subseteq X$, for every $B$ in $sch(\Gamma)$ there exists $Y\to B$ in $FD$ with $Y\subseteq X$. Therefore, in this case, item 3 in Proposition 8 can be stated as: for every $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B)$. Since $\Gamma$ is independent, as mentioned right after Definition 4, this is equivalent ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. The proof is therefore complete. $\Box$
In the following example, we illustrate Proposition 8 and Corollary 1, and we show the impact of our two restrictions on the form of the selection conditions.
@@
In the context of Example 3, we consider the star-table $T$, and its associated m-table $m\_Chase(T)$ as shown in Figure 2, let $Q_{2}$ be the query defined by:
@@
As already noticed after stating Proposition 7, we have ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. On the other hand, the selection condition in this query obviously satisfies our restrictions, since it is independent and involves no key attributes. Applying Proposition 8 or Corollary 1 (because $K_{1}K_{2}\subseteq sch(Q_{2})$) yields the following:
  • The m-tuples $\sigma$ in $m\_Chase(T)$ such that $sch(Q_{i})\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$, are $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ and $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$.
  • Among these m-tuples, only $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ is such that for every $Y\to B$ in $FD$ such that $YB\subseteq X_{i}$, $|{\sf tuples}(\sigma(B))|=1$.
  • This m-tuple satisfies ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
@@
$Q_{3}:$ select $K_{1},K_{2},A^{2}_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ and $M_{1}=m_{1}$) or ($A_{1}^{1}=a^{\prime}_{1}$ and $M_{1}=m^{\prime\prime}_{1}$).
@@
On the other hand, considering the third item in Proposition 8, we emphasize the following: $k_{1}k^{\prime}_{2}a_{2}$ is discarded from the consistent answer because $a_{1}m^{\prime}_{1}$ is in ${\sf tuples}(\sigma_{2}(A_{1}^{1}M_{1}))$ but not in $Sat(\Gamma_{3})$. However, $\sigma_{1}$ can be seen as satisfying the third item in the proposition for the following reasons: the dependencies to consider are $K_{1}\to A_{1}^{1}$ and $K_{1}K_{2}\to M_{1}$, and ${\sf tuples}(\sigma_{1}(A^{1}_{1}))$, respectively ${\sf tuples}(\sigma_{1}(M_{1}))$, is a sub-set of the set of all $A_{1}^{1}$-values, respectively all $M_{1}$-values, occurring in $Sat(\Gamma_{3})$. A characterization of the consistent answer to a query involving a non independent selection condition is currently unknown to the authors.
Now, to illustrate our restriction on the key-attributes with respect to selection conditions, let $Q_{4}:$ select $A_{1}$ from $T$ where ($K_{1}=k_{1}$). It should be clear that the selection condition of $Q_{4}$ is independent, and that ${\sf C\_ans}(Q_{4})=\emptyset$ (because, by Figure 3, neither $a_{1}$ nor $a^{\prime}_{1}$ is associated with $k_{1}$ is every repair). On the other hand, it is easy to see that the conditions in Proposition 8 or Corollary 1 are satisfied by the first two m-tuples of $m\_Chase(T)$ (see Figure 2). A characterization of the consistent answer to a query with a selection condition involving a key attribute is currently unknown to the authors. $\Box$

### v2

@@ 4.2 Repairs of Star-Tables
Let $T$ be a star-table over universe $U$. For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{q}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$. Then, for every $R$ in ${\sf Rep}(T)$, there exists $q_{0}$ in $\{1,\ldots,q\}$ such that $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R})$.
Proof. See Appendix C. $\Box$
@@
We now show how to characterize all repairs of a star-table $T$, based on m_Chase(T)m\_Chase(T). To this end, we define the following three-step process (P):
@@
Step P1. For every $i$ in $\{1,\ldots,n\}$ and every $A_{i}^{j}$ in $sch^{*}(D_{i})$, if $k_{i}$ is a $K_{i}$-value occurring in m_Chase(T)m\_Chase(T), we choose one $A_{i}^{j}$-value in the set $\sigma(A_{i}^{j})$ where $\sigma$ is any m-tuple in m_Chase(T)m\_Chase(T) such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$. Denoting by $\varphi_{i}^{j}(k_{i})$ this value, we notice that $\varphi_{i}^{j}(k_{i})$ is well defined because, thanks to $K_{i}\to A_{i}^{j}$ in $FD$, all $\sigma$ in m_Chase(T)m\_Chase(T) such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$ have the same value over $A_{i}^{j}$.
@@
Step P2. For every $l$ in $\{1,\ldots,p\}$, if $k$ is a $\mathbb{K}$-value occurring in m_Chase(T)m\_Chase(T), we choose one $M_{l}$-value in the set $\sigma(M_{l})$ where $\sigma$ is any m-tuple in m_Chase(T)m\_Chase(T) such that $\mathbb{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$. Denoting by $\varphi_{l}(k)$ this value, we notice that, as above, $\varphi_{l}(k)$ is well defined because, thanks to $\mathbb{K}\to M_{l}$ in $FD$, all $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$ have the same value over $M_{l}$.
@@
Step P3. Denoting by $\varphi$ the set of all $\varphi_{i}^{j}(k_{i})$ and all $\varphi_{l}(k)$ defined above, for every $\sigma$ in m_Chase(T)m\_Chase(T), let $t_{\varphi}(\sigma)$ be the tuple such that $sch(t_{\varphi}(\sigma))=sch(\sigma)$, defined by:
@@
  • If $\mathbb{K}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$, for every $l=1,\ldots,p$ such that $M_{l}\in sch(\sigma)$, let $t_{\varphi}(\sigma).M_{l}=\varphi_{l}(k)$.
@@
In the context of Example 3, $T$ has four repairs as shown in Figure 4. It can be seen from the m-table m_Chase(T)m\_Chase(T) in Figure 2 that, in each of these four tables:
  • the first two rows are generated using the first m-tuple in m_Chase(T)m\_Chase(T), namely
@@
  • the next five rows are generated using the second m-tuple in m_Chase(T)m\_Chase(T), namely
@@
  • the last two rows are generated using respectively the last two m-tuples in m_Chase(T)m\_Chase(T), namely $(k^{\prime}_{1})(k^{\prime\prime}_{2})(a_{1})(m_{1})$ and $(k^{\prime\prime}_{1})(a^{\prime}_{1})(a^{\prime}_{2})$.
@@
  • Step P2. The only measure attribute is here $M_{1}$ and the $\mathbb{K}$-values to be considered are $k_{1}k_{2}$, $k_{1}k^{\prime}_{2}$ and $k^{\prime}_{1}k^{\prime\prime}_{2}$ which are respectively associated with $M_{1}$-values $m_{1}$, $m^{\prime\prime}_{1}$ and $m_{1}$.
  • Step P3. Considering every m-tuple $\sigma$ of the table m_Chase(T)m\_Chase(T) shown in Figure 2, we obtain the following four tuples: $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$ and $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
Hence, $R_{\varphi}$ contains the four tuples above along with the tuples in ${\sf Cons}(\mathcal{T})$ that have been characterized in Example 3. Moreover, inspecting thoroughly the table $R_{1}$ and the set $R_{\varphi}$ shows that the two sets yield the same set of true tuples, corresponding to the same repair of $T$. $\Box$
@@ 5 Consistent Query Answering
    Standard queries (or simply queries), to be defined in the next section. These queries are simple projection-selection queries of the form select $X$ from $T$ where $\Gamma$, where $X$ is a list of attributes and $\Gamma$ an optional selection condition.
@@
    Analytic queries with no group-by clause, defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$, where $X$ and $\Gamma$ are as above, and where aggraggr is an aggregation operator and $M_{i}$ is a measure attribute.
@@
    Analytic queries with a group-by clause, also defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$, where $X$, $\Gamma$, aggraggr and $M_{i}$ are as above.
@@
    Analytic queries with a group-by-having clause, dealt with in Section 6.4. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$, where $X$, $\Gamma$, aggraggr and $M_{i}$ are as above, and where $\Upsilon$ is a boolean expression involving aggregates.
@@ 5.1 Standard Queries
The set of all attributes occurring in $\Gamma$ is called the schema of $\Gamma$, denoted by $sch(\Gamma)$; and the attribute set $X\cup sch(\Gamma)$ is called the schema of $Q$, denoted by $sch(Q)$. Moreover, in what follows, given an attribute set $X$, $\mathcal{T}(X)$ denotes the set of all tuples in $\mathcal{T}$ whose schema is $X$.
@@
Given a selection condition $\Gamma$, the set of tuples satisfying $\Gamma$, denoted by $Sat(\Gamma)$, is inductively defined as follows:
@@
Moreover, based on Proposition 1, it is shown in LauS25 that the two bounds ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ can be easily computed by means of one scan of the m-table m_Chase(T)m\_Chase(T). It turns out that approximating the consistent answer to any query in our approach is polynomial. For example, in the context of Example 3,
@@ 5.2 Consistent Query Answering and Star Schemas
As mentioned in our introductory section, we can compute efficiently exact consistent answers (instead of bounds of exact consistent answers as in LauS25 ) if the selection condition is independent, as stated in the following definition.
@@
For example, let $\Gamma=(A_{1}\leq 0)\vee((A_{1}\geq 4)\wedge(A_{2}=a))$. Clearly, the conjunctive normal form of $\Gamma$ is $((A_{1}\leq 0)\vee(A_{1}\geq 4))\wedge((A_{1}\leq 0)\vee(A_{2}=a))$. Due to the second conjunct, it turns out that $\Gamma$ cannot be written as the conjunction of disjunctions involving only one attribute. Therefore, $\Gamma$ is not independent.
On the other hand, $\Gamma^{\prime}=((A_{1}\leq 4)\vee(A_{1}\geq 0))\wedge(A_{2}=a)$ is obviously independent with $\Gamma^{\prime}(A_{1})=(A_{1}\geq 4)\vee(A_{1}\leq 0)$ and $\Gamma^{\prime}(A_{2})=(A_{2}=a)$, and we have $Sat(\Gamma^{\prime})=\{(a_{1},a_{2})\in adom(A_{1})\times adom(A_{2})~|~(a_{1}\leq 0\vee a_{1}\geq 4)\wedge(a_{2}=a)\}$.
@@
  • From now on, it is assumed, even when not explicitly mentioned, that selection conditions are independent.
@@
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that
@@
    for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$:
    1. (a)
      if $B\in X$, then $|{\sf tuples}(\sigma(B))|=1$,
    2. (b)
      if $B\in sch(\Gamma)$, then ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
Proof. See Appendix E. $\Box$
An important consequence of Proposition 9 is that the time complexity of consistent query answering to projection-selection-join queries on star schemes is quasi-linear (i.e., in $\mathcal{O}(W\cdot\log(W))$, where $W$ is the size of the data warehouse), under the restriction that the selection condition is independent. The following corollary shows that the items in Proposition 9 can be more simply stated when $\mathbb{K}$ is a subset of $X$.
@@
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition. If $\mathbb{K}\subseteq X$, ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that
@@
    $|{\sf tuples}(\sigma(X))|=1$, i.e., ${\sf tuples}(\sigma(X))=\{x\}$,
@@
Proof. The inclusion of $\mathbb{K}$ in $X$ implies that for every $B$ in $sch(Q)\setminus\mathbb{K}$, there exists $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$. Considering moreover that for every $Y\subseteq\mathbb{K}$, $|{\sf tuples}(\sigma(Y))|=1$ always holds, items 1 and 2(a) in Proposition 9 are equivalent to item 2 in the corollary. Similarly, for every attribute $B$ in $sch(\Gamma)$, either $B$ is in $\mathbb{K}$ or there exists $Y\to B$ in $FD$ such that $Y\subseteq\mathbb{K}$. Thus, considering the latter case, item 2(b) in Proposition 9 is equivalent to: for every $B\in sch(\Gamma)\setminus\mathbb{K}$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$. On the other hand if $B$ is in $\mathbb{K}$, we have ${\sf tuples}(\sigma(B))=(b)$ where $b$ is in $Sat(\Gamma(B))$, by item 1 in Proposition 9. Hence item 1 and item 2(b) imply that for every $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B)\subseteq Sat(\Gamma(B))$. Since $\Gamma$ is independent, as mentioned right after Definition 4, this is equivalent ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. The proof is therefore complete. $\Box$
In the following example, we illustrate Proposition 9 and Corollary 1, along with the importance of our restriction on the selection conditions.
@@
In the context of Example 3, we consider the star-table $T$, and its associated m-table m_Chase(T)m\_Chase(T) as shown in Figure 2, let $Q_{2}$ be the query defined by:
@@
As already noticed after stating Proposition 8, we have ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. As the selection condition in this query is obviously independent, applying Proposition 9 or Corollary 1 (because $K_{1}K_{2}\subseteq sch(Q_{2})$) yields the following:
  • The m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) such that $sch(Q_{i})\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$, are $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ and $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$.
  • Among these m-tuples, only $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ is such that: for every $Y\to B$ in $FD$ such that $YB\subseteq X_{i}$, $|{\sf tuples}(\sigma(B))|=1$.
  • The unique m-tuple $\sigma$ of the previous item is such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
@@
$Q_{3}:$ select $K_{1},K_{2},A^{2}_{1}$ from $T$ where
($A_{1}^{1}=a_{1}$ and $M_{1}=m_{1}$) or ($A_{1}^{1}=a^{\prime}_{1}$ and $M_{1}=m^{\prime\prime}_{1}$).
@@
On the other hand, considering the third item in Proposition 9, we emphasize the following: $k_{1}k^{\prime}_{2}a_{2}$ is discarded from the consistent answer because $a_{1}m^{\prime}_{1}$ is in ${\sf tuples}(\sigma_{2}(A_{1}^{1}M_{1}))$ but not in $Sat(\Gamma_{3})$. However, $\sigma_{1}$ can be seen as satisfying the third item in the proposition for the following reasons: the dependencies to consider are $K_{1}\to A_{1}^{1}$ and $K_{1}K_{2}\to M_{1}$, and ${\sf tuples}(\sigma_{1}(A^{1}_{1}))$, respectively ${\sf tuples}(\sigma_{1}(M_{1}))$, is a subset of the set of all $A_{1}^{1}$-values, respectively all $M_{1}$-values, occurring in $Sat(\Gamma_{3})$. A characterization of the consistent answer to a query involving a non independent selection condition is currently unknown to the authors. $\Box$